If f(x)=3x-2 what is (f^-1)(5) and (f^-1)(y)?
Answer by jim_thompson5910(35256)

f(x)=3x-2

y=3x-2

x=3y-2

x+2 = 3y

y = (x+2)/3

(f^-1)(x) = (x+2)/3

Use this to find (f^-1)(5) and (f^-1)(y)